QUESTION M.C.Q (Please answer this question as soon as possible)
You are expecting to receive $100,000 at the end of year 4 and another $100,000 at the end of year 6. What is the present value of the receivables if the interest rate is 5%? $161,005.30
$164,736.38
$154,457.98
$156,891.79
$165,324.49
Answer #1

H299 fi -SUM(H293:H298) D Year Cashflow PVCF PV@5% 0.952380952 0.907029478 0.863837599 0.822702475 0.783526166 0.746215397 PrH299 fc =SUM(H293:H298) 291 Year Cashflow 294 295 PV@5% =1/1.05 =G293/1.05 =G294/1.05 =G295/1.05 =G296/1.05 =G297/1.05 Presen